OpenStudy (anonymous):

divide (2x^3-x^2+3x-1)/(x+2)

OpenStudy (anonymous):

you need more medals for making it look as nice as it does lol

myininaya (myininaya):

lol i think the person left before anyone could answer

myininaya (myininaya):

oh wait maybe he is here i dont know

OpenStudy (anonymous):

\[2 x^2-5 x+13 \text{ with } a \text{ remainder} \text{ of } -27 \]\[\left(\left(2 x^2-5 x+13\right)(x+2)\right)-27 = 2 x^3-x^2+3 x-1 \]

myininaya (myininaya):

2x^2 -5x+13 -------------------------------- x+2 | 2x^3 -x^2 +3x -1 - (2x^3 +4x^2) ----------------------------------- -5x^2 +3x -1 -(-5x^2 -10x) ------------------------- 13x-1 -(13x-26) ---------------- 25

myininaya (myininaya):

thats for correcting my quotient but i believe 25 is the right remainder

OpenStudy (anonymous):

@myininaya Believe all you want Sir. You have made an error in the last multiplication. 13 * (x + 2) is (13x + 26), not (13x - 26) . (13x -1) - (13x + 26) = -27 , the remainder. Refer to the "show steps" in the fourth pane at: http://www.wolframalpha.com/input/?i=divide+%282x^3-x^2%2B3x-1%29%2F%28x%2B2%29

myininaya (myininaya):

yep i made a mistake 13(x+2) is not 13x-26 its 13x+26 ok you are right i didn't see that i'm not a sir

myininaya (myininaya):

also if you just tell me my mistake i will believe you its not what i believe i just thought my math was right so thanks for showing my error and thats all you have to do to make me believe you lol

OpenStudy (anonymous):

Thank you for taking the time to respond. I have posted inadvertent mistakes my self on this math site. Sorry regarding the Sir usage. Perhaps the whole sentence could be viewed as a bit edgy. No was harm intended.